OpenStudy (anonymous):

i need help i will give you a medal For a right triangle that has legs a and b and hypotenuse c, find the length of the side not given here. a=9 c=15

OpenStudy (anonymous):

use \[a ^{2}+b ^{2}=c ^{2}\] you are given a and c

OpenStudy (anonymous):

9^2+b^2=15^2 right

OpenStudy (anonymous):

yup then solve for b

OpenStudy (anonymous):

ok i will give you a medal right now

OpenStudy (anonymous):

cool thanks :)

OpenStudy (anonymous):

your welcome can we be friends

OpenStudy (anonymous):

suree np

OpenStudy (anonymous):

so the answer is 12 correct or i am worng

OpenStudy (anonymous):

hold up a sec 81+b^2=225 -81 -81 b^2=114 sqrt.b=sqrt114 b=12

OpenStudy (anonymous):

am i right

OpenStudy (anonymous):

yes I got 12 as well

OpenStudy (anonymous):

k i poseted another one

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

k